OpenStudy (anonymous):

Select the approximate values of x that are solutions to f(x) = 0, where f(x) = -5x2 + 7x + 2. {–0.24, 1.64} {-5, 7} {–1.40, –0.40} {–2.50, 3.50}

OpenStudy (anonymous):

what number could you put in for x that would make the output zero

OpenStudy (anonymous):

since it is multiple choice you could just plug them into the function

OpenStudy (anonymous):

i dont understand the question sorry thats why i asked for some help? can someone please help me with this for i have no idea

myininaya (myininaya):

you can see which of the pairs get you output closest to 0 are you can use the quadratic formula and take the approximations form the exact solutions

myininaya (myininaya):

\[ax^2+bx+c=0 \text{ implies } x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

myininaya (myininaya):

So first step identify a,b,c